OpenStudy (anonymous):

factor the polynomial 12y^2-27

Parth (parthkohli):

It's not apolynomial. I think you didn't complete the question. You've typed a binomial.

OpenStudy (anonymous):

Difference of squares....

OpenStudy (callisto):

12y^2-27 = 3 (4x^2 - 9) = 3 (2x-3)(2x+3)

OpenStudy (anonymous):

it said "factor the polynomail ,if possible. if the polynomail cannot be factored , write the Pime

OpenStudy (anonymous):

Binomial is a polynomial

Parth (parthkohli):

^ Actually no, there can be monomials, binomials, trinomials and polynomials. Trinomial is included in polynomials, not binomials.